the difference between two numbers is 15. four times the smaller number plus twice the larger number is 120. what are the two numbers?
Answer by whenindoubt(13)

Given information:
x= smaller number
y= larger number
y-x=15
therefore y= x+15


4x + 2y = 120
4x + 2(x+15) = 120
4x + 2x + 30 = 120
6x + 30 = 120
6x = 90
x= 15
therefore


x = 15
y = 30